The force of resistance on an object moving through a fluid is proportional to the
square of its velocity. The initial velocity of an object of mass 1 kg moving through the
fluid is 40 ms−1
. After 5 s the velocity reduces to 20 ms−1
. Calculate its velocity after
10 s.
1
Expert's answer
2019-03-15T12:43:54-0400

Write Newton's second law in differential form:


"m\frac{dv}{dt}=-kv^2,"

integrate:


"-\frac{1}{v}=-\frac{k}{m}t+C."

Substituting v=40 m/s and t=0 find "C":


"C=-\frac{1}{40}\text{ s/m}."

Substituting v=20 m/s and t=5 s find "k":


"k=0.005\text{ kg/m}."

Substituting t=10 s determine what is required:


"v=13.3\text{ m/s}."